OpenStudy (anonymous):

What is the speed of a rocket that travels 8000 meters in 12 seconds?

OpenStudy (anonymous):

Find how much it travels in a second: 8000/12 = 666.666 m/s (or 2000/3 as a simple fraction) Multiply by 3.6 666.666 * 3.6 = 2400 km/h

OpenStudy (anonymous):

If you notice you could do it directly as well. \[\frac{ 8000 }{ 12 }\times3.6=2400_{km/h}\]

OpenStudy (mrnood):

it doesn't specify any units speed can be mph, kph, m/s, yards per fortnight. a perfectly good answer would be 8000m per 12 sec

OpenStudy (mrnood):

If you wish to stick to sensible 'standard' units then kph is not the preferred SI unit You should calculate the speed in m/s 8000/12 = 666.666 m/s